OpenStudy (nicole143):

@mathmale Hello, can you help me with solving this equation? -2/x+4 = 4/x+3

OpenStudy (nicole143):

If you can help me with the first bit of getting the LCD, I think I could do it from there (: Or I may be wrong ha

OpenStudy (anonymous):

\[\frac{ -2 }{x+4 }=\frac{ 4 }{ x+3 } \] Am i correct ?

OpenStudy (nicole143):

Yes

OpenStudy (anonymous):

cross-multiply

OpenStudy (nicole143):

Example?

OpenStudy (anonymous):

4x+16=-2x-6 4x+2x=-6-16 6x=-22 x=?

OpenStudy (nicole143):

Oh, -11/3 ?

OpenStudy (anonymous):

\[\frac{ a }{b+c }=\frac{ d }{ e+f },ae+af=db+dc\]

OpenStudy (anonymous):

correct .Have you followed?

OpenStudy (nicole143):

Yes, thank you. I think I got it(:

OpenStudy (anonymous):

yw